# Understood: Real-Time Communication Support for Adults with ADHD Using Mixed Reality

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06submissionid: 8878ccs: Human-centered computing Accessibility systems and tools
@@ 1. Introduction
Attention-Deficit Hyperactivity Disorder (ADHD) is a persistent neurodevelopmental disorder marked by patterns of inattention, hyperactivity, and impulsivity that significantly impair daily functioning (kooij2010european). Epidemiological studies estimate that ADHD affects $3\%$-$10\%$ of school-aged children and $1\%$-$6\%$ of adults worldwide, with approximately one to two thirds of childhood cases continuing into adulthood, necessitating lifelong adaptation strategies (wender2001adults). While diagnostic criteria have traditionally focused on the core behavioral symptoms in children (katsarou2024identifying; mathers2006aspects; martinussen2006working; mikami2015social; storebo2019social; wilkes2017pragmatic; wenchen2016language), recent research increasingly highlights the characteristics and impact of ADHD in adults (surman2013understanding; safren2006cognitive; solanto2011cognitive; wender2001adults; barkley2007adhd-ch13). Specifically, ADHD in adults presents distinct phenotypes: whereas hyperactivity is more salient in childhood and often facilitates earlier diagnosis, adulthood manifestations tend to shift toward emotional dysregulation and executive dysfunction, reflecting more complex and internalized symptoms. (katzman2017adult). Notably, adults with ADHD experience marked impairments in executive functions (boonstra2005executive; spencer2008triple), working memory (ossmann2003inhibition), and cognitive flexibility (roshani2020comparison), which together contribute to the broader difficulties associated with the disorder. These neural and behavioral differences collectively contribute to dysfunction across multiple life domains.
@@
The impact of these differences becomes particularly evident in interpersonal communication (westby2021adhd). Clinical studies demonstrate that many individuals with ADHD report persistent social communication challenges (wenchen2016language; theodoratou2024communication; mathers2006aspects), despite these difficulties not being included in the core diagnostic criteria for the disorder (american2013diagnostic). These communication difficulties are rooted in the neurobiological and behavioral characteristics of ADHD and manifest in distinct patterns across various interaction contexts. Clinical evaluations indicate that individuals with ADHD often experience a phenomenon known as “stuckness” during conversations, where they struggle to maintain a fluid dialogue (katsarou2024identifying; tannock2018adhd). Additionally, many exhibit impaired discourse cohesion, a result of working memory deficits that lead to tangential narratives and difficulties in resolving references during conversations (baker1992attention). Furthermore, a significant portion display impulse-driven communication breakdowns, where premature responses disrupt conversational turn-taking norms (bellani2011language; mathers2006aspects). Collectively, these challenges create a “conversational paradox” — individuals with ADHD often possess strong verbal intelligence, yet their executive dysfunction impedes effective message delivery.
@@
Current non-medicinal interventions have shown limited efficacy in addressing the real-time communication challenges inherent in ADHD management. While evidence-based behavioral interventions (collins2005strategy; reimers1987acceptability), including Cognitive Behavioral Therapy (CBT) (solanto2011cognitive; safren2006cognitive) , Social Skills Training (SST) (mikami2017social; storebo2019social; mikami2015social), and Self-Regulation Training (barkley2004attention; surman2013understanding) exhibit clinical effectiveness in controlled settings through systematic skill acquisition, their real-world applicability remains constrained. SST, in particular, offers promise in improving social interaction deficits by fostering structured behavioral modifications and enhancing interpersonal relationships. However, the skills acquired through SST often fail to generalize beyond simulated scenarios, revealing a critical gap between therapeutic training and the complexities of real-world social interactions (willis2019stand). Similarly, while CBT effectively targets maladaptive thought patterns, it reliance on cognitive flexibility and self-monitoring limits its applicability for adults with ADHD, who often experience executive dysfunction that hampers real-time conversational adaptability (antshel2014cognitive). Self-regulation training, which incorporates positive thinking and metacognitive strategies, enhances awareness of emotional states but does not sufficiently bridge the gap between internal regulation and external appropriate social responses during spontaneous interactions (cibrian2022potential). Meanwhile, technology-integrated supports present potential solutions. Virtual Reality (VR) systems provide immersive, controlled environments for graded exposure to challenging interactions (wong2023effectiveness). However, VR often fosters passive, prompted learning, which may hinder patients from independently identifying and addressing interpersonal challenges. Similarly, wearable biofeedback devices leverage neural activity modulation to improve cognitive and emotional regulation, enabling individuals to adjust their brain activity through training (swingle2008biofeedback; doren2019sustained). While these devices show promise in mitigating ADHD-related dysfunctions, concerns remain regarding their intrusive nature and the lack of regulatory oversight, which often leads to inconsistent application and unreliable outcomes (geppert2019neuro).
@@
The above landscape reveals three key gaps: G1: Sustainable transfer of therapeutic outcomes - current systems rely on retrospective analysis, restricting skill-building to clinical contexts and making it difficult to maintain behavioral improvements in natural social interactions; G2: Empowerment of user autonomy - existing implementations treat patients as passive diagnostic subjects, rather than empowering them to actively regulate their own communicative patterns; and G3: Minimally intrusive systems - many solutions disrupt natural conversation flow with overt notifications. Our work seeks to bridge these gaps by integrating computational linguistics, wearable devices, and neurodiversity-informed design.
@@
Therefore, our study seeks to explore the following four research questions (RQ1–RQ4). To begin, we aim to understand communication challenges faced by individuals with ADHD and identify suitable support strategies. This leads us to two foundational research questions: RQ1: What communication challenges do individuals with ADHD encounter, and what are the current intervention strategies? and RQ2: What strategies could better support individuals with ADHD in mitigating these communication challenges? To investigate RQ1, we conducted two semi-structured interview studies involving four licensed clinicians and 12 adults with ADHD. These interviews revealed three primary barriers: (1) deficits in working memory; (2) disfluency and pauses in conversation; and (3) attentional anchoring failure (e.g., abrupt topic shifts. In addressing RQ2, we carried out a design study involving clinicians, researchers, and patients, that began with an exploratory design brainstorming session to generate diverse concepts targeting these communication issues. We then developed storyboards to visualize the concepts and held a design workshop to further refine the supportive approaches. MR emerged as a preferred medium due to its potential to deliver information non-intrusively, minimize conversational disruptions, and support polite interaction.
@@
Based on feedback from clinicians, researchers, and patients, we defined five design goals and developed Understood, an MR system implemented on Microsoft HoloLens 2111https://learn.microsoft.com/en-us/hololens/, designed to support adults with ADHD in real-world communication. Understood incorporates three core features: (1) real-time conversation summarization to address working memory limitations and reduce cognitive load, (2) context-aware subsequent word suggestions explicitly activated by users during moments of disfluency or hesitation, and (3) topic shift detection and reminding to help mitigate off-topic transitions.
@@
In this study, we further explored two key research questions concerning the Understood system: RQ3: How effectively does Understood support communication? and RQ4: What is the usability of Understood? To answer these questions, we conducted a within-subjects study in which participants completed communication tasks under two conditions: (1) using Understood and (2) without it. We collected both quantitative performance metrics and qualitative feedback through semi-structured interviews, enabling a comprehensive evaluation of the system’s communication support (RQ3) and usability (RQ4). The findings consistently demonstrated that Understood significantly improved communication fluency for individuals with ADHD while maintaining a high level of usability. In summary, our contributions are as follows:
@@ 2.1. ADHD and Communication Challenges
ADHD is a neurodevelopmental condition characterized by persistent patterns of inattention, impulsivity, and hyperactivity, which disrupt various aspects of daily life (wender2001adults; furman2005attention). These symptoms are particularly challenging in conversational settings, where individuals with ADHD often face difficulties with both verbal and non-verbal communication. These challenges arise from the core symptoms of the condition, manifesting in distinctive ways during interactions and hindering their ability to engage effectively in interpersonal communication (wenchen2016language; theodoratou2024communication).
One significant challenge observed in individuals with ADHD is a phenomenon known as “Stuckness” during conversations (katsarou2024identifying). This occurs when a person becomes mentally fixed on a specific topic, word, or thought, making it difficult to transition smoothly or adapt to the natural flow of dialogue (tannock2018adhd). As a result, conversations may feature long pauses, repeated phrases, or abrupt topic shifts, often leaving conversational partners confused or disengaged. Another communication barrier is incoherent logic in discourse (baker1992attention). Individuals with ADHD often struggle to organize their thoughts, leading to fragmented or tangential speech patterns. They might jump between unrelated ideas or fail to provide sufficient context, making it difficult for others to grasp their intended message. This disorganization is closely linked to impairments in working memory and executive functioning – cognitive processes responsible for planning, organizing, and regulating thoughts  (martinussen2006working; mcinnes2003listening). Additionally, inappropriate wording and poor self-regulation in speech are common, with impulsivity often leading to interruptions, blurting out responses, or using contextually inappropriate language (bellani2011language; mathers2006aspects). These behaviors can cause misunderstandings, disrupt conversational flow, and create social friction. Together, these challenges present significant obstacles to effective communication in both personal and professional scenarios.
Given that many of these difficulties stem from deficits in executive functions, such as working memory, cognitive flexibility, and inhibitory control (bellani2011language), our study aims to enhance conversational fluency, cohesion, and overall social interactions for individuals with ADHD by addressing these specific communication challenges. To achieve this, we propose a system that offers real-time feedback and guidance, supporting individuals with ADHD during their conversations.
@@ 2.2. Behavioral Interventions for Communication Challenges in ADHD
Effectively addressing communication challenges in individuals with ADHD requires interventions that enhance fluency and coherence during conversations. The conventional non-medical approach primarily relies on behavioral interventions (collins2005strategy; reimers1987acceptability), which focus on developing foundational skills through structured training (rajeh2017interventions; daley2014behavioral; caye2019treatment). These interventions typically target areas such as active listening (clayton2021behavioral), staying on topic (anderson1991improving), and self-regulating speech (graham1992developing).
Behavioral interventions have traditionally been the primary approach to supporting individuals with ADHD in improving communication skills. These interventions focus on developing strategies to manage impulsivity, enhance attention, and organize thoughts during conversations (theodoratou2024communication; wilkes2017pragmatic). One key method is CBT (solanto2011cognitive; safren2006cognitive), which aims to alter maladaptive thought patterns that disrupt communication. By utilizing self-monitoring and cognitive restructuring, individuals learn to identify and manage behaviors such as distractibility and impulsivity. Antshel et al. (antshel2014cognitive) found that CBT can effectively improve conversational skills and self-regulation in individuals with ADHD, although its success is contingent on the individual’s engagement and consistent application of strategies. Another widely-used method is SST (mikami2017social; storebo2019social; mikami2015social), which targets specific conversational competencies such as turn-taking, initiating and maintaining dialogue, and interpreting social cues. Through role-playing exercises, individuals practice real-world scenarios to improve their social interactions. Research by Willis et al. (willis2019stand) suggests that SST can reduce social isolation and improve communication, though it may not address cognitive deficits like working memory, which are essential for spontaneous conversation. Additionally, Self-regulation training (barkley2004attention; surman2013understanding) incorporates techniques like mindfulness to help individuals control impulsive speech and emotional responses during conversations. Studies, such as those by Cibrian et al. (cibrian2022potential), show that self-regulation training can improve attention and emotional responses in ADHD, particularly in challenging social situations. However, this approach requires significant time and effort to achieve lasting results and may be less effective if ADHD symptoms remain unmanaged.
@@ 2.3. Technology-Integrated Supports for Communication Challenges
Researchers are increasingly developing technology-integrated supports that use computational methods to analyze communication difficulties in ADHD and provide interactive support (wong2023effectiveness; raeisian2025confident; Sehlin2018experiences; keshav2019digital). These emerging technologies serve as valuable extensions to conventional interventions by offering real-time feedback and dynamic assistance during conversations. For example, Virtual Reality (VR) technologies offer immersive environments for practicing communication in realistic settings. Wong et al. (wong2023effectiveness) developed a VR-based system designed to help individuals with ADHD practice social skills. While these systems provide a safe, controlled environment for practice, their practical implementation remains limited due to the constrained range of scenarios they provide. Some of these solutions combine wearable devices with biofeedback, or specifically neurofeedback systems to monitor physiological indicators, such as heart rate, skin conductance, and brain activity (swingle2008biofeedback; doren2019sustained; geppert2019neuro). Swingle et al. (swingle2008biofeedback) demonstrated that, through training, individuals can learn to modify their brainwave patterns to better manage cognitive challenges such as inattention and impulsivity, thereby indirectly improving communication skills. However, despite demonstrated effectiveness, the application of these methods in clinical settings often deviates from research-based protocols due to insufficient regulation, leading to inconsistent outcomes and reduced reliability.
@@
In addition to research focused on ADHD, several studies have explored communication support for individuals with various communication challenges (zulfikar2024memoro; bermejo2020vimes; hohenstein2023artificial; ou2024academic; elsahar2019augmentative; yu2024holoaac; curtis2024looking). Elsahar et al. (elsahar2019augmentative) provided a comprehensive review of tools designed to assist individuals with speech and language impairments, evaluating them in terms of accessibility, affordability, complexity, portability, and typical conversation speed. Similarly, Hohenstein et al. (hohenstein2023artificial) and Ou et al. (ou2024academic) investigated how artificial intelligence shapes users’ perceptions and influences linguistic and social interactions in both everyday and academic settings. Regarding specific tools, Curtis et al. (curtis2024looking) evaluated three high-fidelity prototypes—embody projection, worn audio devices, and headset-driven MR – through the lens of visual modalities, highlighting the potential of current and emerging MR technologies in enhancing communication. Zulfikar et al. (zulfikar2024memoro), on the other hand, focused on auditory modalities and propose a wearable memory-augmentation system. Although these works do not specifically target individuals with ADHD, they offer valuable design insights that can inform our own implementation.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025copyright: ccconference: The 38th Annual ACM Symposium on User Interface Software and Technology; September 28-October 1, 2025; Busan, Republic of Koreabooktitle: The 38th Annual ACM Symposium on User Interface Software and Technology (UIST ’25), September 28-October 1, 2025, Busan, Republic of Koreadoi: 10.1145/3746059.3747602isbn: 979-8-4007-2037-6/2025/09submissionid: 8878ccs: Human-centered computing Accessibility systems and tools
@@
[…]
Three-panel illustration labeled A, B, and C showing a scenario involving communication challenges and assistive technology. (1) Panel A: Depicts several individuals experiencing communication difficulties. One person has a thought bubble labeled “(BLANK),” another says “um…” while looking confused, and a third person shows jumbled thoughts involving unrelated icons like a flower and a soccer ball. (2) Panel B: Shows a person wearing a HoloLens headset conversing with another person. Icons on the left indicate system features: conversation summarization, subsequent word suggestion, and topic shifting detection. (3) Panel C: The same two individuals converse smoothly, both smiling with matching flower icons in their speech bubbles, indicating improved mutual understanding.
@@ 1. Introduction
[…]
@@
The impact of these differences becomes particularly evident in interpersonal communication (Westby and Watson, 2021). Clinical studies demonstrate that many individuals with ADHD report persistent social communication challenges (Helland et al., 2016; Theodoratou, 2024; Mathers, 2006), despite these difficulties not being included in the core diagnostic criteria (Association et al., 2013). These communication difficulties are rooted in the neurobiological and behavioral characteristics of ADHD and manifest in distinct patterns across various interaction contexts. Clinical evaluations indicate that individuals with ADHD often experience “Stuckness” during conversations, where they struggle to maintain a fluid dialogue (Katsarou et al., 2024; Tannock, 2018), exhibit impaired discourse cohesion with tangential narratives (Baker and Cantwell, 1992), and disrupt conversational turn-taking with premature responses (Bellani et al., 2011; Mathers, 2006). Collectively, these challenges create a “conversational paradox” — individuals with ADHD often possess strong verbal intelligence, yet their executive dysfunction impedes effective message delivery.
@@
Current non-medicinal interventions have shown limited efficacy in addressing the real-time communication challenges inherent in ADHD management. While evidence-based behavioral interventions (Collins et al., 2005; Reimers et al., 1987), including Cognitive Behavioral Therapy (CBT) (Solanto, 2011; Safren, 2006), Social Skills Training (SST) (Mikami et al., 2017; Storebø et al., 2019; Mikami, 2015), and Self-Regulation Training (Barkley, 2004; Surman et al., 2013) exhibit clinical effectiveness in controlled settings, their applicability in real-world communication remains constrained. For instance, SST promotes structured behavioral modifications, but acquired skills often fail to generalize beyond simulated scenarios (Willis et al., 2019). CBT targets maladaptive thought patterns, yet its reliance on cognitive flexibility limits its real-time adaptability (Antshel and Olszewski, 2014). Similarly, self-regulation training enhances awareness of emotional states but does not sufficiently support spontaneous, socially appropriate responses (Cibrian et al., 2022). Technology-integrated supports present potential solutions: VR systems provide immersive, controlled exposure (Wong and Qin, 2023), but often foster passive, prompted learning; Wearable biofeedback devices aim to improve regulation through brain activity training (Swingle, 2008; Doren et al., 2019), though concerns remain regarding their intrusiveness and the lack of regulatory oversight, which may leads to unreliable outcomes (Enriquez-Geppert et al., 2019). Additionally, prior work in HCI, mainly with children, has demonstrated autonomy-supportive system designs for ADHD (Stefanidi et al., 2023; Silva et al., 2023; Sonne et al., 2016), but these remain limited in addressing real-time communication needs in adults.
@@
The above landscape reveals three key gaps: G1: Sustainable transfer of therapeutic outcomes - current systems rely on retrospective analysis, restricting skill-building to clinical contexts and making it difficult to maintain behavioral improvements in natural social interactions; G2: Empowerment of user autonomy in ADHD adults - Although much of the prior work, particularly with children, has supported my in system-mediated interventions for ADHD, existing implementations often still treat ADHD adults as passive diagnostic subjects, rather than empowering them to actively regulate their own communicative patterns; and G3: System with minimal user disruption - many existing solutions disrupt the user’s natural conversational flow through overt notifications and cognitively demanding interactions. Our work seeks to bridge these gaps by integrating computational linguistics, wearable devices, and neurodiversity-informed design.
@@
Therefore, our study seeks to explore the following four research questions (RQ1–RQ4). To begin, we aim to understand communication challenges faced by individuals with ADHD and identify suitable support strategies. This leads us to two foundational research questions: RQ1: What communication challenges do individuals with ADHD encounter, and what are the current intervention strategies? and RQ2: What strategies could better support individuals with ADHD in mitigating these communication challenges? To investigate RQ1, we conducted two semi-structured interview studies involving four licensed clinicians and 12 adults with ADHD. These interviews revealed three primary barriers: (1) deficits in working memory; (2) disfluency and pauses in conversation; and (3) attentional anchoring failure (e.g., abrupt topic shifts). In addressing RQ2, we carried out a design study involving clinicians, researchers, and ADHD people, that began with an exploratory design brainstorming session to generate diverse concepts targeting these communication issues. We then developed storyboards to visualize the concepts and held a design workshop to further refine the supportive approaches. MR emerged as a preferred medium due to its potential to deliver information non-intrusively, minimize conversational disruptions, and support polite interaction.
@@
Based on feedback from clinicians, researchers, and ADHD people, we defined five design goals and developed Understood, an MR system implemented on Microsoft HoloLens 2111https://learn.microsoft.com/en-us/hololens/, designed to support adults with ADHD in real-world communication. Understood incorporates three core features: (1) real-time conversation summarization to address working memory limitations and reduce cognitive load, (2) context-aware subsequent word suggestions explicitly activated by users during moments of disfluency or hesitation, and (3) topic shift detection and reminding to help mitigate off-topic transitions.
@@
In this study, we further explored two key research questions concerning the Understood system: RQ3: How effectively does Understood support communication? and RQ4: What is the usability of Understood? To answer these questions, we conducted a within-subjects study in which participants completed communication tasks under two conditions: (1) using Understood and (2) without it. We collected both quantitative performance metrics and qualitative feedback through semi-structured interviews, enabling a comprehensive evaluation of the system’s communication support (RQ3) and usability (RQ4). The findings consistently demonstrated that when evaluated under a within-subject, lab-based setting, Understood significantly improved communication fluency for individuals with ADHD while maintaining a high level of usability. In summary, our contributions are as follows:
@@ 2.1. ADHD and Communication Challenges
ADHD is a neurodevelopmental condition characterized by persistent patterns of inattention, impulsivity, and hyperactivity, which disrupt various aspects of daily life (Wender et al., 2001; Furman, 2005). These symptoms are particularly challenging in conversational settings, where individuals with ADHD often face difficulties with both verbal and non-verbal communication. These challenges arise from the core symptoms of the condition, manifesting in distinctive ways during interactions and hindering their ability to engage effectively in interpersonal communication (Helland et al., 2016; Theodoratou, 2024).
One significant challenge observed in individuals with ADHD is a phenomenon known as “Stuckness” during conversations (Katsarou et al., 2024). This occurs when a person becomes mentally fixed on a specific topic, word, or thought, making it difficult to transition smoothly or adapt to the natural flow of dialogue (Tannock, 2018). As a result, conversations may feature long pauses, repeated phrases, or abrupt topic shifts, often leaving conversational partners confused or disengaged. Another communication barrier is incoherent logic in discourse (Baker and Cantwell, 1992). Individuals with ADHD often struggle to organize their thoughts, leading to fragmented or tangential speech patterns. They might jump between unrelated ideas or fail to provide sufficient context, making it difficult for others to grasp their intended message. This disorganization is closely linked to impairments in working memory and executive functioning – cognitive processes responsible for planning, organizing, and regulating thoughts  (Martinussen and Tannock, 2006; McInnes et al., 2003). Additionally, inappropriate wording and poor self-regulation in speech are common, with impulsivity often leading to interruptions, blurting out responses, or using contextually inappropriate language (Bellani et al., 2011; Mathers, 2006). These behaviors can cause misunderstandings, disrupt conversational flow, and create social friction. Together, these challenges present significant obstacles to effective communication in both personal and professional scenarios.
Given that many of these difficulties stem from deficits in executive functions, such as working memory, cognitive flexibility, and inhibitory control (Bellani et al., 2011), our study aims to enhance conversational fluency, cohesion, and overall social interactions for individuals with ADHD by addressing these specific communication challenges. To achieve this, we propose a system that offers real-time feedback and guidance, supporting individuals with ADHD during their conversations.
@@
[…]
A flowchart diagram showing the formative study procedure, with blue boxes representing research activities and gray boxes representing research entities. The process starts with two groups under ”Interviews”: ”Experts” (illustrated as medical professionals) and ”ADHD Individuals” (illustrated as people appearing distressed). Arrows lead to identified ”Challenges (Ch1–Ch3)” and initial ”Findings (F1–F2)”. These feed into a ”Design Study” phase, including: ”Design Brainstorming” (group discussion), ”Design Concepts” (a person with a lightbulb and gears), ”Storyboards” (a person wearing a HoloLens), and a ”Design Workshop” (group activity). The workshop produces more ”Findings (F3–F10),” leading to final ”Design Goals (DG1–DG5),” depicted as a person confidently pointing forward.
@@ 2.2. Behavioral Interventions for Communication Challenges in ADHD
Effectively addressing communication challenges in individuals with ADHD requires interventions that enhance fluency and coherence during conversations. The conventional non-medical approach primarily relies on behavioral interventions (Collins et al., 2005; Reimers et al., 1987), which focus on developing foundational skills through structured training (Rajeh et al., 2017; Daley et al., 2014; Caye et al., 2019). These interventions typically target areas such as active listening (Clayton et al., 2021), staying on topic (Anderson and Sharpe, 1991), and self-regulating speech (Graham et al., 1992).
Behavioral interventions have traditionally been the primary approach to supporting individuals with ADHD in improving communication skills. These interventions focus on developing strategies to manage impulsivity, enhance attention, and organize thoughts during conversations (Theodoratou, 2024; Wilkes-Gillan et al., 2017). One key method is CBT (Solanto, 2011; Safren, 2006), which aims to alter maladaptive thought patterns that disrupt communication. By utilizing self-monitoring and cognitive restructuring, individuals learn to identify and manage behaviors such as distractibility and impulsivity. Antshel et al. (Antshel and Olszewski, 2014) found that CBT can effectively improve conversational skills and self-regulation in individuals with ADHD, although its success is contingent on the individual’s engagement and consistent application of strategies, it also requires cognitive flexibility and real-time self-monitoring, which are often impaired in adults with ADHD. Another widely-used method is SST (Mikami et al., 2017; Storebø et al., 2019; Mikami, 2015), which targets specific conversational competencies such as turn-taking, initiating and maintaining dialogue, and interpreting social cues. Through role-playing exercises, individuals practice real-world scenarios to improve their social interactions. Research by Willis et al. (Willis et al., 2019) suggests that SST can reduce social isolation and improve communication, though its benefits may remain confined to structured sessions and fail to generalize to spontaneous, real-world conversations, where working memory and cognitive flexibility are crucial. Additionally, Self-regulation training (Barkley, 2004; Surman et al., 2013) incorporates techniques like mindfulness to help individuals control impulsive speech and emotional responses during conversations. Studies, such as those by Cibrian et al. (Cibrian et al., 2022), show that self-regulation training can improve attention and emotional responses in ADHD, particularly in challenging social situations. However, while it enhances emotional awareness, this does not necessarily translate into timely or socially appropriate responses during real-time interactions. Moreover, this approach requires significant time and effort to achieve lasting results and may be less effective if ADHD symptoms remain unmanaged.
@@ 2.3. Technology-Integrated Supports for Communication Challenges
Researchers are increasingly developing technology-integrated supports that use computational methods to analyze communication difficulties in ADHD and provide interactive support (Wong and Qin, 2023; Raeisian Parvari et al., 2025; Sehlin et al., 2018; Keshav et al., 2019). These emerging technologies serve as valuable extensions to conventional interventions by offering real-time feedback and dynamic assistance during conversations. For example, Virtual Reality (VR) technologies offer immersive environments for practicing communication in realistic settings. Wong et al. (Wong and Qin, 2023) developed a VR-based system designed to help individuals with ADHD practice social skills. While these systems provide a safe, controlled environment for practice, their practical implementation remains limited due to the constrained range of scenarios they provide. Some of these solutions combine wearable devices with biofeedback, or specifically neurofeedback systems to monitor physiological indicators, such as heart rate, skin conductance, and brain activity (Swingle, 2008; Doren et al., 2019; Enriquez-Geppert et al., 2019). Swingle et al. (Swingle, 2008) demonstrated that, through training, individuals can learn to modify their brainwave patterns to better manage cognitive challenges such as inattention and impulsivity, thereby indirectly improving communication skills. However, despite demonstrated effectiveness, the application of these methods in clinical settings often deviates from research-based protocols due to insufficient regulation, leading to inconsistent outcomes and reduced reliability. Some research has also explored how system design can support autonomy in individuals with ADHD (Stefanidi et al., 2023; Silva et al., 2023; Sonne et al., 2016). Stefanidi et al. (Stefanidi et al., 2023) proposed approaches that move beyond symptom-oriented interventions by engaging the broader care ecosystem. However, these efforts have primarily targeted children, with limited attention to autonomy-supportive communication tools for adults with ADHD.
@@
In addition to research focused on ADHD, several studies have explored communication support for individuals with various communication challenges (Zulfikar et al., 2024; Bermejo et al., 2020; Hohenstein et al., 2023; Ou et al., 2024; Elsahar et al., 2019; Yu et al., 2024; Curtis et al., 2024). Elsahar et al. (Elsahar et al., 2019) provided a comprehensive review of tools designed to assist individuals with speech and language impairments, evaluating them in terms of accessibility, affordability, complexity, portability, and typical conversation speed. Similarly, Hohenstein et al. (Hohenstein et al., 2023) and Ou et al. (Ou et al., 2024) investigated how artificial intelligence shapes users’ perceptions and influences linguistic and social interactions in both everyday and academic settings. Regarding specific tools, Curtis et al. (Curtis et al., 2024) evaluated three high-fidelity prototypes—embody projection, worn audio devices, and headset-driven MR – through the lens of visual modalities, highlighting the potential of current and emerging MR technologies in enhancing communication. Zulfikar et al. (Zulfikar et al., 2024), on the other hand, focused on auditory modalities and propose a wearable memory-augmentation system. Although these works do not specifically target individuals with ADHD, they offer valuable design insights that can inform our own implementation.